OpenStudy (anonymous):

Addition Method, I just want to make sure I am doing this right before I move on: a+b=3 & 3a+b=13

OpenStudy (anonymous):

-3(a+b=3) 1(3a+b=13) -3a-3b=-9 3a+b=13 end up with -3b = -9 b =13 -2b = 4 divide b = -2 then plug in -2 to one of orig. problems 3a-2+2=13+2 3a=15 divide a=5 The order would be? ( 5,-2)

OpenStudy (anonymous):

check... it must solve both original equations.

Directrix (directrix):

>>The order would be? ( 5,-2) That is correct. @Jlynn89